Two capacitors C1 =18.0 mF and C2 = 36.0 mF are connected in series and a 12.0V battery is connected across the two capacitors. Find
(i) the equivalent capacitance and the energy stored in this equivalent capacitance.
(ii) Find the energy stored in each individual capacitor
1
Expert's answer
2021-02-17T11:10:25-0500

"C_1 = 18.0mkF" "C_2 = 36.0mkF" "U = 12.0V"

Two capacitors are connected in series.

(i) the equivalent capacitance and the energy stored in this equivalent capacitance.

the equivalent capacitance for connected in series is "C = \large\frac{C_1*C_2}{C_1+C_2}= \frac{18*36}{18+36}" "=12mkF"

 the energy stored in this equivalent capacitance. "W = \large\frac{CU^2}{2}" "= \large\frac{12*144}{2}" "=864mkJ"

(ii) Find the energy stored in each individual capacitor

the charges in capacitors and total charge for system is equal to each other in connected series.

"q = q_1 = q_2 = CU= 12mkF*12V = 144mkC"

for the the first capacitor "W_1 = \large\frac{q_1^2}{2C_1}" "= \large\frac{144*10^{-6}*144*10^{-6}}{2*18*10^{-6}}" "= 576mkJ"

for the the second capacitor "W_2 = \large\frac{q_2^2}{2C_2}" "= \large\frac{144*10^{-6}*144*10^{-6}}{2*36*10^{-6}} = 288mkJ"